OpenStudy (anonymous):

What is the simplified form of 2 over x squared minus x minus 1 over x ? x minus 1 over the quantity x times x plus 1 1 minus x over the quantity x times x plus 1 3 minus x over the quantity x times x minus 1 x plus 2 over the quantity x times x minus 1

OpenStudy (johnweldon1993):

\[\frac{ 2 }{ x^2 - x } - \frac{ 1 }{ x }\] Right?

OpenStudy (johnweldon1993):

Well assuming it IS that....notice how the 1st denominator can be factored... \[\frac{ 2 }{ x(x - 1) } - \frac{ 1 }{ x }\] Notice now that all we have to do...is multiply the second fraction by (x - 1) to have a common denominator... So this would come out to ... \[\frac{ 2 }{ x(x - 1) } - \frac{ 1 }{ x } \times \frac{(x - 1)}{(x - 1)}\] Simplified down a bit and we have \[\frac{ 2 }{ x(x - 1) } - \frac{ x - 1 }{ x(x - 1) }\] We can now put these over the common denominator \[\frac{ 2 - (x - 1) }{ x(x - 1) }\] Lets simplify the top a bit...distribute the '-' sign... \[\frac{ 2 - x + 1 }{ x(x - 1) }\] And finally we arrive at... \[\frac{ 3 - x }{ x(x - 1) }\] Which answer choice is that?